Deoro Company has identified the following overhead activities, costs, and activity drivers for the coming year: Activity Expected Cost Activity Driver Activity Capacity Setting up equipment $469,800 Number of setups 580 Ordering costs 348,000 Number of orders 17,400 Machine costs 747,000 Machine hours 41,500 Receiving 462,000 Receiving hours 10,500 Deoro produces two models of dishwashers with the following expected prime costs and activity demands: Model A Model B Direct materials $647,000 $779,000 Direct labor $471,000 $484,000 Units completed 15,600 7,200 Direct labor hours 6,000 2,400 Number of setups 380 200 Number of orders 6,200 11,200 Machine hours 23,900 17,600 Receiving hours 4,000 6,500 The company's normal activity is 8,400 direct labor hours. Required: 1. Determine the unit cost for each model using direct labor hours to apply overhead. Round intermediate calculations and final answers to nearest cent. 2. Determine the unit cost for each model using the four activity drivers. Round your answers to nearest cent.
Answer #1

Overhead Rate=Total Expected cost / Total direct labor hours

=(469800+348000+747000+462000)/(6000+2400)

=241.2857 or 241

Cost per unit = (Direct materials + direct labor )/units completed + direct labor hours * overhead rate /units completed)

Model A cost per unit =(647000+471000)/15600+ 6000*241/15600

=1118000/15600+ 1446000/15600

= 71.67+ 92.69=164

Model B= (779000+484000) /7200 + 2400*241/7200

=1263000/7200 + 578400/7200

=175.42 + 80.33= 256

B. Unit cost =[ (direct materials + Direct labor ) + (cost of setup equipment +ordering cost machine cost +. Receiving cost )]/ units completed

Model A = ((647000+471000)+ 469800× 380/580 + 348000* 6200/17400 +747000*23900/41500 +462000*4000/10500 )/15600

=( 647000+471000+ 307800+ 124000+ 430200+176000 )/15600

= 2156000/15600= 138

Model B :( (779000+484000) +469800*200/580+ 348000*11200/17400+ 747000*17600/41500 +462000*6500/10500))/7200

= (779000+484000+162000+ 224000+ 316800+ 286000)/7200

= 2251800/7200= 313